Ticket Number: HEA-NOC/19981105-02 Ticket Status: UPDATE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Avoca.heanet.ie
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19981110 18:05 UTC
Ticket Closed: Problem End: 19981110 19:28 UTC
Upgrades are being carried out on Avoca.heanet.ie (a.k.a.
www.heanet.ie, ftp.heanet.ie, cache.heanet.ie, listserv.heanet.ie)
Tuesday's upgrade was only partially successful, due to problems with
the equipment shipped that only became obvious during the upgrade.
The memory was still installed, and a map of Avoca's SCSI layout was
taken which will speed the RAID installation still to come.
These services were unavailable for the duration of the outage:
National Information Server website
National web cache
The upgrade should ease pressure on the system and result in better
performance for LISTSERV and the proxy cache in particular.
LISTSERV mail has been queued and is now clearing.
Digital engineer installed RAM and noted SCSI layout. A further outage
will be scheduled & this ticket updated. The outage is expected to
take place later this week.
Time to Fix:
System, most services: 1h 23m
NIS and associated websites: 1h 33m